A mass mkg is suspended by a weightless string of length l a horizontal force is applied to displace it slowly until the string makes 30degree with the initial vertical direction what is the work done by applied force?

Grade:10

1 Answers

Diwankar singh
6 years ago
Mass=m,length=l. Work done=Force*distance. Force=mass*acceleration =m*g. --------(1) Cos30°=b/h where,h=length of string=l and b=displacement=? Cos30°=b/l Cos30°=√3/2Therefore, √3/2=b/lSo,b=l√3/2=displacement. ----------(2)From (1) and (2)Work done=mgl√3/2. Ans.